OpenStudy (anonymous):

solve the inequality and graph the solution x+1 2/5 < 6 8/10 (greater than or equal to sign)

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@misty1212 hiii

OpenStudy (anonymous):

\[x+1\frac{ 2 }{ 5 }<6\frac{ 8 }{ 10 }\] \[x<6\frac{ 4 }{ 5 }-1\frac{ 2 }{ 5 }\] \[x<\frac{ 34 }{ 5 }-\frac{ 7 }{ 5 }\] or \[x<\frac{ 27 }{ 5 }\] \[x <5\frac{ 2 }{ 5 }\]